OpenStudy (anonymous):

Derivative of this function.? Could someone please show and explain step by step how to take this derivative (2x^3 -3)^5 over 10

OpenStudy (anonymous):

I need the anti derivative sorry

OpenStudy (anonymous):

\[\int\frac{(2x^3-3)^5}{10}~dx\] It'll be tedious, but expanding the binomial is the way to go: \[(2x^3-3)^5=32x^{15}-240x^{12}+720x^{9}-1080x^{6}+810x^{3}-243\] \[\frac{1}{10}\int\left(32x^{15}-240x^{12}+720x^{9}-1080x^{6}+810x^{3}-243\right)~dx\] If there's a simpler way, I'm not seeing it right away.